OpenStudy (anonymous):

use the chain rule to find the derivative: y=1/cot^5(x-2)

OpenStudy (anonymous):

the right format: \[y=\frac{ 1 }{ \cot⁵(x⁻²) }\]

OpenStudy (anonymous):

so I can get functions f(x)=1/x, g(x)=cot⁵(x⁻²), h(x)=x⁻²

OpenStudy (anonymous):

so dy/dx=f'(g(h(x)))*g'(h(x))*h'(x)

OpenStudy (anonymous):

uh, what's the derivative of cot⁵? is the derivative of cot x=-csc²x, then is it, i dunno, -csc¹⁰x?

OpenStudy (zehanz):

Before differentiating, simplify: \(f(x)=\dfrac{1}{\cot^5(x-2)}=\tan^5(x-2)\). This is because cotx=cosx/sinx and tanx =sinx/cosx. Now you need to know the derivative of tanx. You have two options: \[(\tan x)'= \tan^2 x+1 = \frac{1}{\cos^2 x}\] (I don't think you are asked to prove this in this problem, so you can use this). Remember: \(\tan^5(x-2)=(\tan(x-2))^5\), so if you need the derivative, you must first differentiate the 5th power and then multiply with the derivative of tan(x-2): \(5\tan^4(x-2) \cdot (\tan (x-2))'=...\)

OpenStudy (anonymous):

sorry there was a typo in the first post. It should have been\[y=\frac{ 1 }{ \cot⁵(x⁻²) }\] So should I pull another function out of the original? as in\[dy/dx=5*(\tan(x⁻²))⁴*\frac{ 1 }{ \cos²(x⁻²) }*\frac{ -2 }{ x³ }\]

OpenStudy (zehanz):

Yes, but the derivative of \(x^{-2}\) is \(-2x^{-3}=-\dfrac{2}{x^3}\), so the derivative of your function is \(\dfrac{dy}{dx}=5\tan^4(x^{-2})\cdot \dfrac{1}{\cos^2(x^{-2})}\cdot-\dfrac{2}{x^3}\). I will try to make it look a little friendlier: \(\dfrac{dy}{dx}=-\dfrac{10\tan^4(x^{-2})}{x^3\cos^2(x^{-2})}\). Still not pretty, but a little shorter :)

OpenStudy (anonymous):

thanks @zehans

OpenStudy (zehanz):

YW!